# MMP-A*: Multimodal Perception Enhanced Incremental Heuristic Search on Path Planning

### v1

@@ 1 Introduction
Path planning is a core problem in robotics and autonomous navigation, requiring the computation of an optimal or near-optimal route from start to goal while avoiding obstacles, and serving as a foundational capability for mobile robots, autonomous vehicles, industrial automation, and virtual environments where navigation efficiency directly impacts safety and scalability hart1968formal; abd2015comprehensive; gonzalez2015review. Classical search-based methods, particularly A* and its derivatives, are widely utilized for their theoretical guarantees of completeness and optimality under admissible heuristics koenig2004lifelong; karaman2011sampling. However, increasing scale and clutter trigger exponential computational costs in these algorithms. While specialized A* variants offer optimizations, they often rely on specific environmental priors, rendering them brittle and unable to generalize to the arbitrary, unstructured geometries of real-world settings.
Large language models (LLMs) have increasingly been applied across a wide range of domains, including robotics, where they support tasks such as high-level task planning and action selection tariq2025robust; doma2024llm; joublin2024copal. In particular, LLM-A* meng2024llm frameworks employ LLMs to generate intermediate waypoints that guide the search process toward semantically meaningful regions, significantly reducing expansion cost compared to vanilla A*. While this integration introduces promising global reasoning, the text-only modality of LLMs cannot encode fine-grained geometric or topological structures caglar2024can; wei2025plangenllms; cao2025large. Consequently, in complex environments with dense or irregular barriers, LLM-generated waypoints often become redundant, misplaced, or geometrically infeasible, leading to unstable heuristic evaluations and degraded performance, sometimes worse than A* itself. Vision-Language Models (VLMs) partially address this issue by incorporating spatial grounding from visual inputs, enabling recognition of navigable areas and geometric relationships ye2025vlm. However, inherent limitations in long-horizon planning render both LLMs and VLMs ineffective as standalone end-to-end motion planners aghzal2023can; aghzal2024look; yang2025guiding. This constraint necessitates a paradigm shift: rather than serving as autonomous controllers, these models may be more effectively utilized as supportive modules within a structured planning framework. The complementary strengths of LLMs and VLMs, balancing high-level semantic reasoning with precise geometric grounding, thus motivate a multimodal integration designed to unify abstract cognition with physical spatial constraints.
@@ 2 Related Work
Pathfinding has long been a fundamental problem in artificial intelligence, robotics, and computer graphics. The A* algorithm hart1968a remains foundational for combining heuristic estimation with optimality guarantees, inspiring numerous extensions to improve efficiency and adaptability. Notable A* variants typically align with two distinct operational paradigms. In known environments, strategies such as IDA* korf1985depth, RTA*/LRTA* korf1990real, and SMA* russell1992memory address the equilibrium between computational cost, memory constraints, and responsiveness. In contrast, dynamic environments necessitate algorithms like D* stentz1994optimal, LPA* koenig2004lifelong or Bug1 lumelsky1987path which are explicitly engineered to facilitate efficient incremental replanning. To enhance scalability, hierarchical and structure-aware approaches like HPA* botea2004near and JPS harabor2011online further reduce search overhead, underscoring the enduring influence of heuristic search in modern path planning. In this context, our work specifically targets large-scale, known environments, aiming to address the scalability challenges posed by high-complexity static maps.
@@
LLMs have recently demonstrated impressive reasoning and generalization abilities across natural language processing and decision-making domains naveed2023comprehensive; chang2024survey. Their capacity to decompose instructions and infer latent goals has motivated applications in high-level planning and embodied control erdogan2025plan; yang2025planllm; zhao2023large. To address the computational complexity of path planning, works like meng2024llm and tariq2025robust utilize LLMs to generate high-level waypoints, effectively pruning the search space and guiding geometric planners toward optimal solutions in large-scale environments. Conversely, some approaches wang2024llmˆ; doma2024llm; oelerich2024language focus on trajectory feasibility, employing LLMs to translate semantic instructions into dynamic cost map updates or task-space constraints, ensuring the generation of safe, collision-free paths under complex restrictions. However, LLMs still exhibit notable limitations in long-horizon planning and spatial reasoning, especially in tasks requiring continuous geometric understanding aghzal2023can; ilharco2020probing; patel2021mapping; abdou2021can. To address this, we propose a hybrid framework that boosts high-level planning by coordinating LLMs with VLMs to anchor reasoning in physical space, coupled with a low-level A* planner to deliver mathematically accurate pathfinding results.
@@
VLMs have emerged as a promising solution to these spatial reasoning limitations by jointly processing visual and linguistic information zhang2024vision; han2025multimodal; xu2025llava. Through visual grounding, VLMs directly perceive environmental geometry, allowing them to assess navigability and spatial patterns that remain opaque to text-only models. Recent frameworks utilize VLMs to ground search in visual reality, employing techniques like semantically-biased sampling ye2025vlm or visual affordance prompting chen2025affordances to steer planners toward safe, goal-relevant regions. Beyond generation, research also explores using VLMs as zero-shot reward models aghzal2025evaluating, evaluating trajectory compliance against complex constraints. Despite their perceptual capabilities, VLMs struggle to translate pixel-level information into continuous geometric coordinates, making visual input alone inadequate for rigorous motion planning baghaei2025follow; colan2025assessing.
Algorithm 1 MMP-A* Algorithm with Adaptive Decay for Path Planning
1:Require: START state $s_{0}$, GOAL state $s_{g}$, OBSTACLE set $\mathcal{S}_{obs}$, heuristic function $h()$, cost function $g()$, Large Language Model $llm()$, Vision Language Model $vlm()$, adaptive decay factor $\alpha\in(0,1)$
2:Initialize OPEN list $\mathcal{O}_{\text{open}}\leftarrow\{s_{0}\}$, CLOSE list $\mathcal{C}_{\text{close}}\leftarrow\{\}$, TARGET list $\mathcal{T}\leftarrow vlm(llm(s_{0},s_{g},\mathcal{O}))$, TARGET state $t\leftarrow\mathcal{T}_{\text{start}}$, $g(s_{0})\leftarrow 0$, $f(s_{0})\leftarrow h(s_{0})$, $k\leftarrow 0$
@@ 3.1 Problem Formulation
To solve the formulation above, we establish the classical A* algorithm as the foundational planner. Unlike blind search methods, A* directs the exploration towards $s_{g}$ by minimizing a heuristic-guided cost function $f(s)$ for every visited node $s$:
@@ 3.3 LLM-Guided Waypoint Generation
Despite these structural safeguards, the LLM operates on abstract textual data and lacks intrinsic spatial grounding. Consequently, the generated waypoints may exhibit geometric inaccuracies. Critically, as the environment scales or obstacle density increases, the textual representation becomes unwieldy, leading to prohibitive computational costs and context-length bottlenecks. These limitations necessitate the subsequent VLM-based refinement stage, which employs visual perception to robustly validate and optimize the waypoint candidates.
@@ 3.4 VLM-Based Visual Refinement
To mitigate spatial hallucinations from the text-based phase, we introduce a VLM refinement stage that validates the coarse list $T$. The model is prompted with two aligned visual inputs: (i) a raw occupancy grid marking barriers and $\{s_{0},s_{g}\}$, and (ii) a visualization layer overlaying LLM-proposed waypoints onto the grid.
Our prompt engineering operationalizes the concept of a ”valid checkpoint” as a node that maintains safety margins from walls and avoids congestion. The refinement workflow proceeds as follows:
@@ 3.5 Adaptive Heuristic Search Integration
$h_{LLM\!A^{*}}(n)=h_{A^{*}}(n)+cost(n,t)$ (2)
becomes unstable. Consequently, the search tends to diverge toward suboptimal areas, greatly increasing computation and slowing convergence. This limitation arises from the static treatment of LLM-generated waypoints, where the algorithm maintains excessive confidence in each waypoint regardless of its reliability.
To address this issue, we introduce an adaptive decay factor $\alpha$ to dynamically reduce the influence of LLM-generated waypoints over time. Specifically, the heuristic function is reformulated as
$h_{MMP\!A^{*}}(n)=h_{A^{*}}(n)+\alpha^{k}cost(n,t),$ (3)
where $cost(n,t)$ measures the cost from the current node $n$ to the current waypoint $t$, $h_{A^{*}}(n)$ is the cost from $n$ to the goal, and $\alpha\in(0,1)$ decays exponentially with each waypoint switch ($k$ denotes the index of the current waypoint).
@@ 4.1 Dataset
    Environmental Complexity is modulated via obstacle density, culminating in Level 5, a rigorous benchmark featuring 12 intertwined barrier clusters and deep dead-ends designed to challenge multimodal reasoning.
@@
For a detailed specification of these parameters and generation protocols, please refer to Appendix D.
@@ 4.2 Experimental Setup
Our evaluation focuses on efficiency and scalability by benchmarking MMP-A* against A* and LLM-A*. We assess computational cost and memory usage via Operation and Storage Ratios, computed as the geometric mean of performance ratios between MMP-A* and the A* baseline ($\frac{\text{MMP-A*}}{\text{A*}}$), where lower values indicate superior resource efficiency. Path optimality is evaluated via Relative Path Length, while system reliability is measured by the Valid Path Ratio, representing the proportion of successfully generated collision-free trajectories. Comprehensive results are summarized in Table 1.
@@ 4.3 Experimental Results
The comprehensive evaluation in Table 12 reveals that purely generative models fail to capture topological constraints of complex environments, yielding Valid Path Ratios below $10\%$ and trajectories inflated by up to $30\%$ due to geometric hallucinations. While integrating language guidance with search via LLM-A* resolves the validity issue, it inadvertently incurs a substantial computational penalty, with operation ratios surging to $191.6\%$ in some configurations as the search algorithm is forced to explore misleading regions suggested by noisy heuristics. In stark contrast, our proposed MMP-A* achieves a superior balance of efficiency and optimality across all model backbones. By leveraging VLM-based visual pruning to eliminate infeasible waypoints before the search commences and employing an adaptive heuristic decay, MMP-A* not only guarantees $100\%$ validity but also drastically reduces resource consumption, bringing operation and storage costs down to $81.0\%$ and $76.0\%$ respectively while maintaining path lengths within $2\%$ of the optimal solution.
@@ 4.4 Ablation Study
As shown in Table 11 in Appendix and Figures 4, incorporating adaptive decay slightly increases the operation and storage ratios, by about 5–10%, since the planner performs extra expansions to refine its heuristic. Nevertheless, it remains markedly more efficient than vanilla A*, while consistently producing smoother and shorter paths. This small overhead is offset by a 1–2% reduction in relative path length, confirming that adaptive decay effectively stabilizes the search by gradually lowering reliance on uncertain early waypoints. When varying the decay coefficient $\alpha$, we observe a clear efficiency–optimality trade-off: higher $\alpha$ values reduce operation cost but yield longer paths due to stronger waypoint bias, whereas lower $\alpha$ values encourage broader exploration and shorter trajectories at slightly higher cost (Figures 2). Overall, moderate $\alpha$ values achieve the best balance, preserving efficiency gains while maintaining near-optimal path quality.
@@
Figure 2: Alpha-Decay Sensitivity Analysis. Operation ratio (bars) and relative path length (lines) of LLM-A* and MMP-A* under varying decay coefficients $\alpha$. Each subplot corresponds to a different LLM–VLM pairing.
@@
We evaluated three prompting schemes to analyze their influence on waypoint generation and search efficiency. As summarized in Table 10 and Figure 5c, CoT prompts generally yield better reasoning consistency but incur higher computational cost, with operation ratios often 10–20% above Few-Shot setups. RePE introduces recursive self-evaluation, which improves local waypoint precision and slightly reduces path length (about 0.5–1.0%), though at the expense of additional memory usage. Among these, Few-Shot remains the most lightweight and stable baseline, while RePE offers the best trade-off between cost and accuracy when integrated with MMP-A*.
Figure 4: Impact of Adaptive Decay on Efficiency and Path Quality. Comparison between LLM-A* and MMP-A* with and without the adaptive decay mechanism. The plots show changes in operation, storage, and relative path length ratios (lower is better).
Figure 5: Overall visualization of key experiments. (a) Complex Environment Experiment: Operation and storage ratios across increasing map complexity levels. (b) Scale Robustness Experiment: Growth factors of operations and storages across varying map sizes. (c) Comparison of Prompt Engineering Strategy

### v2

@@ Abstract
Code is available at: https://github.com/langkhachhoha/MMP-ASTAR
**footnotetext: Equal contribution.
@@ 1 Introduction
Path planning is a core problem in robotics and autonomous navigation, requiring the computation of an optimal or near-optimal route from start to goal while avoiding obstacles, and serving as a foundational capability for mobile robots, autonomous vehicles, industrial automation, and virtual environments where navigation efficiency directly impacts safety and scalability Hart et al. (1968a); Abd Algfoor et al. (2015); González et al. (2015). Classical search-based methods, particularly A* and its derivatives, are widely utilized for their theoretical guarantees of completeness and optimality under admissible heuristics Koenig et al. (2004); Karaman and Frazzoli (2011). However, increasing scale and clutter trigger exponential computational costs in these algorithms. While specialized A* variants offer optimizations, they often rely on specific environmental priors, rendering them brittle and unable to generalize to the arbitrary, unstructured geometries of real-world settings.
Large Language Models (LLMs) have increasingly been applied across a wide range of domains, including robotics, where they support tasks such as high-level task planning and action selection Tariq et al. (2025); Doma et al. (2024); Joublin et al. (2024). In particular, LLM-A* Meng et al. (2024) frameworks employ LLMs to generate intermediate waypoints that guide the search process toward semantically meaningful regions, significantly reducing expansion cost compared to vanilla A*. While this integration introduces promising global reasoning, the text-only modality of LLMs cannot encode fine-grained geometric or topological structures Caglar et al. (2024); Wei et al. (2025); Cao et al. (2025). Consequently, in complex environments with dense or irregular barriers, LLM-generated waypoints often become redundant, misplaced, or geometrically infeasible, leading to unstable heuristic evaluations and degraded performance, sometimes worse than A* itself. Vision-Language Models (VLMs) partially address this issue by incorporating spatial grounding from visual inputs, enabling recognition of navigable areas and geometric relationships Ye et al. (2025). However, inherent limitations in long-horizon planning render both LLMs and VLMs ineffective as standalone end-to-end motion planners Aghzal et al. (2023, 2024); Yang et al. (2025b). This constraint necessitates a paradigm shift: rather than serving as autonomous controllers, these models may be more effectively utilized as supportive modules within a structured planning framework. The complementary strengths of LLMs and VLMs, balancing high-level semantic reasoning with precise geometric grounding, thus motivate a multimodal integration designed to unify abstract cognition with physical spatial constraints.
@@ 2 Related Work
Pathfinding has long been a fundamental problem in artificial intelligence, robotics, and computer graphics. The A* algorithm Hart et al. (1968b) remains foundational for combining heuristic estimation with optimality guarantees, inspiring numerous extensions to improve efficiency and adaptability. Notable A* variants typically align with two distinct operational paradigms. In known environments, strategies such as IDA* Korf (1985), RTA*/LRTA* Korf (1990), and SMA* Russell (1992) address the equilibrium between computational cost, memory constraints, and responsiveness. In contrast, dynamic environments necessitate algorithms like D* Stentz (1994), LPA* Koenig et al. (2004) or Bug1 Lumelsky and Stepanov (1987) which are explicitly engineered to facilitate efficient incremental replanning. To enhance scalability, hierarchical and structure-aware approaches like HPA* Botea et al. (2004) and JPS Harabor and Grastien (2011) further reduce search overhead, underscoring the enduring influence of heuristic search in modern path planning. In this context, our work specifically targets large-scale, known environments, aiming to address the scalability challenges posed by high-complexity static maps.
@@
LLMs have recently demonstrated impressive reasoning and generalization abilities across natural language processing and decision-making domains Naveed et al. (2023); Chang et al. (2024); Ha et al. (2025). Their capacity to decompose instructions and infer latent goals has motivated applications in high-level planning and embodied control Erdogan et al. (2025); Yang et al. (2025a); Zhao et al. (2023). To address the computational complexity of path planning, works like Meng et al. (2024) and Tariq et al. (2025) utilize LLMs to generate high-level waypoints, effectively pruning the search space and guiding geometric planners toward optimal solutions in large-scale environments. Conversely, some approaches wang2024llmˆ; Doma et al. (2024); Oelerich et al. (2024) focus on trajectory feasibility, employing LLMs to translate semantic instructions into dynamic cost map updates or task-space constraints, ensuring the generation of safe, collision-free paths under complex restrictions. However, LLMs still exhibit notable limitations in long-horizon planning and spatial reasoning, especially in tasks requiring continuous geometric understanding Aghzal et al. (2023); Ilharco et al. (2020); Patel and Pavlick (2021); Abdou et al. (2021). To address this, we propose a hybrid framework that boosts high-level planning by coordinating LLMs with VLMs to anchor reasoning in physical space, coupled with a low-level A* planner to deliver mathematically accurate pathfinding results.
@@
VLMs have emerged as a promising solution to these spatial reasoning limitations by jointly processing visual and linguistic information Zhang et al. (2024); Han et al. (2025); Xu et al. (2025). Through visual grounding, VLMs directly perceive environmental geometry, allowing them to assess navigability and spatial patterns that remain opaque to text-only models. Recent frameworks utilize VLMs to ground search in visual reality, employing techniques like semantically-biased sampling Ye et al. (2025) or visual affordance prompting Chen et al. (2025) to steer planners toward safe, goal-relevant regions. Beyond generation, research also explores using VLMs as zero-shot reward models Aghzal et al. (2025), evaluating trajectory compliance against complex constraints. Despite their perceptual capabilities, VLMs struggle to translate pixel-level information into continuous geometric coordinates, making visual input alone inadequate for rigorous motion planning Baghaei et al. (2025); Colan et al. (2025).
Algorithm 1 MMP-A* Algorithm with Adaptive Decay
1:Require: START state $s_{0}$, GOAL state $s_{g}$, OBSTACLE set $\mathcal{S}_{obs}$, heuristic function $h()$, cost function $g()$, $llm()$, $vlm()$, adaptive decay factor $\alpha\in(0,1)$
[…]
@@ 3.1 Problem Formulation
To solve the formulation, we establish the classical A* algorithm as the foundational planner. Unlike blind search methods, A* directs the exploration by minimizing a heuristic-guided cost function $f(s)$ for every visited node $s$:
@@ 3.3 LLM-Guided Waypoint Generation
Despite these structural safeguards, the LLM operates on abstract textual data and lacks intrinsic spatial grounding. Consequently, the generated waypoints may exhibit geometric inaccuracies. Critically, as the environment scales or obstacle density increases, the textual representation becomes unwieldy, leading to prohibitive computational costs and context-length bottlenecks. These limitations necessitate the subsequent stage, which employs visual perception to robustly validate and optimize the waypoint candidates.
@@ 3.4 VLM-Based Visual Refinement
To mitigate spatial hallucinations from the text-based phase, we introduce a VLM refinement stage that validates the coarse list $T$. The model is prompted with two aligned visual inputs: (i) a raw occupancy grid marking barriers and $\{s_{0},s_{g}\}$, and (ii) a visualization layer overlaying LLM-proposed waypoints onto the grid. Our prompt engineering operationalizes the concept of a valid checkpoint as a node that maintains safety margins from walls and avoids congestion. The refinement workflow proceeds as follows:
@@
We prioritize visual verification over direct generation to exploit the resolution independence of the top-down view. Unlike coordinate-based generation, which struggles with the combinatorial explosion of large grids, a visual approach allows us to abstract the map into a fixed-size image. This ensures that the VLM’s ability to interpret global topology remains robust regardless of the map’s actual scale.
@@ 3.5 Adaptive Heuristic Search Integration
$h_{LLM-\!A^{*}}(n)=h_{A^{*}}(n)+cost(n,t_{k})$ (2)
becomes unstable. Here $cost(n,t_{k})$ measures the cost from the current node $n$ to the current waypoint $t_{k}$. Consequently, the search tends to diverge toward suboptimal areas, greatly increasing computation and slowing convergence. This limitation arises from the static treatment of LLM-generated waypoints, where the algorithm maintains excessive confidence in each waypoint regardless of its reliability. To address this issue, we introduce an adaptive decay factor $\alpha$ to dynamically reduce the influence of LLM-generated waypoints over time. Specifically, the heuristic function is reformulated as
$h_{MMP-\!A^{*}}(n)=h_{A^{*}}(n)+\alpha^{k}cost(n,t_{k}),$ (3)
where $\alpha\in(0,1)$ decays exponentially with each waypoint switch ($k$ denotes the index of the current waypoint).
@@
3.6 Theoretical Analysis
We analyze the theoretical properties of the adaptive decay heuristic employed in MMP-A*.
Proposition 3.1 (Bounded Suboptimality).
Let $h_{A^{*}}(n)$ be an admissible heuristic and define
$h_{\text{MMP-A}^{*}}(n)=h_{A^{*}}(n)+\alpha^{k}\cdot c(n,t_{k}),\quad\alpha\in(0,1),$
where $t_{k}$ denotes the current waypoint and $c(n,t_{k})$ is the estimated cost from node $n$ to $t_{k}$. Assume that there exists a finite constant $D_{\max}$ such that $c(n,t_{k})\leq D_{\max}$ for all $n$ and $k$, e.g., the maximum shortest-path distance in the finite grid. Then, for any finite $k$,
$h_{\text{MMP-A}^{*}}(n)\leq h^{*}(n)+\alpha^{k}D_{\max},$
where $h^{*}(n)$ denotes the true optimal cost-to-go. Consequently, the heuristic exhibits bounded additive suboptimality with an error term $\alpha^{k}D_{\max}$.
Proof Sketch.
By admissibility of $h_{A^{*}}$, we have $h_{A^{*}}(n)\leq h^{*}(n)$. Since the waypoint-dependent term $c(n,t_{k})$ is bounded by $D_{\max}$ and scaled by $\alpha^{k}$, the heuristic introduces a bounded additive overestimation that decays exponentially with the waypoint index. $\square$
Proposition 3.2 (Pointwise Convergence).
For any node $n$,
$\lim_{k\to\infty}h_{\text{MMP-A}^{*}}(n)=h_{A^{*}}(n).$
Proof Sketch.
Since $\alpha\in(0,1)$, the decay factor satisfies $\lim_{k\to\infty}\alpha^{k}=0$. As a result, the waypoint guidance term vanishes asymptotically, and the heuristic converges pointwise to the admissible base heuristic $h_{A^{*}}$. $\square$
Proposition 3.3 (Robustness against Long-Horizon Spatial Hallucinations).
In large-scale environments where the sequence length $k$ is significant, VLM reliability typically degrades, leading to high-error waypoints (hallucinations). Under the LLM-A* formulation ($\alpha=1$), a misleading waypoint $t_{k}$ imposes a persistent heuristic penalty $c(n,t_{k})$, forcing the planner to detour towards the erroneous location regardless of the path length. In contrast, MMP-A* attenuates this risk via the decay factor $\alpha^{k}$. For large $k$, the influence of the misleading waypoint vanishes:
$\lim_{k\to\infty}\alpha^{k}\cdot c(n,t_{k})=0.$
This property ensures that as the problem scale increases, MMP-A* automatically decouples from potentially unreliable VLM guidance and reverts to the admissible goal-directed behavior of $h_{A^{*}}(n)$, effectively bypassing misleading traps that would otherwise entrap a LLM-A* planner.
@@ 4.1 Dataset
    Environmental Complexity is modulated via obstacle density, culminating in Level 5, a rigorous benchmark featuring approximate 12 intertwined barrier clusters and deep dead-ends designed to challenge multimodal reasoning.
@@
Figure 3 illustrates samples from our curated datasets. For a detailed specification of these parameters and generation protocols, please refer to Appendix D.
@@ 4.2 Experimental Setup
Our evaluation focuses on efficiency and scalability by benchmarking MMP-A* against A* and LLM-A*. We assess computational cost and memory usage via Operation and Storage Ratios, computed as the geometric mean of performance ratios between MMP-A* and the A* baseline ($\frac{\text{MMP-A*}}{\text{A*}}$), where lower values indicate superior resource efficiency. Path optimality is evaluated via Relative Path Length, while system reliability is measured by the Valid Path Ratio, representing the proportion of successfully generated collision-free trajectories. Comprehensive results are summarized in Table 1. Figure 8 in Appendix visualizes representative sample results, demonstrating how MMP-A* allows the planner to bypass misleading waypoints and yield smoother trajectories.
@@ 4.3 Experimental Results
The comprehensive evaluation in Table 12 reveals that purely generative models fail to capture topological constraints of complex environments, yielding Valid Path Ratios below $10\%$ and trajectories inflated by up to $30\%$ due to geometric hallucinations. While integrating language guidance with search via LLM-A* resolves the validity issue, it inadvertently incurs a substantial computational penalty, with operation ratios surging to $191.6\%$ in some configurations as the search algorithm is forced to explore misleading regions suggested by noisy heuristics. In stark contrast, our proposed MMP-A* achieves a superior balance of efficiency and optimality across all model backbones. By leveraging VLM-based visual pruning to eliminate infeasible waypoints before the search commences and employing an adaptive heuristic decay, MMP-A* not only guarantees $100\%$ validity but also drastically reduces resource consumption, bringing operation and storage costs down to $81.0\%$ and $76.0\%$ respectively while maintaining path lengths within $2\%$ of the optimal solution. We discuss the comparative efficiency of LLM–VLM models and their limitations in Appendix  E.
@@ 4.4 Ablation Study
As shown in Table 11 in Appendix and Figure 2, incorporating adaptive decay slightly increases the operation and storage ratios, by about 5–10%, since the planner performs extra expansions to refine its heuristic. Nevertheless, it remains markedly more efficient than vanilla A*, while consistently producing smoother and shorter paths. This small overhead is offset by a 1–2% reduction in relative path length, confirming that adaptive decay effectively stabilizes the search by gradually lowering reliance on uncertain early waypoints. When varying the decay coefficient $\alpha$, we observe a clear efficiency–optimality trade-off: higher $\alpha$ values reduce operation cost but yield longer paths due to stronger waypoint bias, whereas lower $\alpha$ values encourage broader exploration and shorter trajectories at slightly higher cost. Overall, moderate $\alpha$ values achieve the best balance, preserving efficiency gains while maintaining near-optimal path quality.
@@
Figure 2: Alpha-Decay Sensitivity Analysis. Operation ratio (bars) and relative path length (lines) of LLM-A* and MMP-A* under varying decay coefficients $\alpha$.
@@
We evaluated three prompting schemes to analyze their influence on waypoint generation and search efficiency. As summarized in Table 10 and Figure 10 in Appendix, CoT prompts generally yield better reasoning consistency but incur higher computational cost, with operation ratios often 10–20% above Few-Shot setups. RePE introduces recursive self-evaluation, which improves local waypoint precision and slightly reduces path length (about 0.5–1.0%), though at the expense of additional memory usage. Among these, Few-Shot remains the most lightweight and stable baseline, while RePE offers the best trade-off between cost and accuracy when integrated with MMP-A*.
Figure 4: (a) Complex Environment Experiment: Operation and storage ratios across increasing map complexity levels. (b) Scale Robustness Experiment: Growth factors of operations and storages across varying map sizes.